OpenStudy (anonymous):

Naomi says she is going to study 3 hours the first day this week and then is going to add on thirty minutes each day for the next seven days. Write a recursive formula to represent Naomi's study pledge.

OpenStudy (anonymous):

at first tell me what you think about this question, to how we solve it ?

OpenStudy (anonymous):

I think I figured out the recursive formula. But I also realized I need and explicit formula.

OpenStudy (anonymous):

yes, but first you need to look the question back again.

OpenStudy (anonymous):

The recusive is\[A(n)=a _{n-1} + 0.5\]

OpenStudy (anonymous):

the question said naomi study 3 hours daily, that means, if we take a function which discribe this, so we can write this as\[F(x)=30\]

OpenStudy (anonymous):

It said only on the first day does she study 3 hrs

OpenStudy (anonymous):

The second day she studies 3.5 hrs and so on

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

so is the formula I wrote correct

OpenStudy (anonymous):

but you say's that also she increase time by (1/2)hrs, per day

OpenStudy (anonymous):

how

OpenStudy (anonymous):

i did it in minutes.

OpenStudy (anonymous):

so its \[A(n)=180+30^{n-1}\] sorry i wrote it wrong last time

OpenStudy (anonymous):

So do I need to write them both in the same terms?

OpenStudy (anonymous):

like for the recursive \[A(n)=a _{n-1}+30\]

OpenStudy (anonymous):

ok,but if she study 3hrs first day, and increase the time of study 1/2hrs per day, so nex day she study for additional (1/2)hrs, and one another next it will increase (1/2)hrs, more

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

So 3hrs on day 1, 3.5 on day 2, 4 on day 3...

OpenStudy (anonymous):

thanks for the help.

OpenStudy (anonymous):

so, i can write to describe this as like a function, as\[f(x)=3+(1/2)x\]where f(x)=total time she study, and x is day after 1st day she has started study.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

bt wouldn't it be \[f(x)=3+(1/2)^{x}\]

OpenStudy (anonymous):

and besides i need a formula not a function, or are they the same.

OpenStudy (anonymous):

the formula that i think is look like that\[f(x)=3+0.5x\] this will give you correct total time on that day,she study for, like 1st day she study for 3 hrs, so 2nd day she study for \[1+(1/2)hrs\]

OpenStudy (anonymous):

sorry,\[3+(1/2)\]

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

Thanks